OpenStudy (anonymous):

can anyone help me with this problem use log b 3=1.099 and /or log b 9=2.197 to find log b 27 the b are below the log

OpenStudy (anonymous):

log b 27=3*(log b 3)=3*1.099=answer; here * means multiply.....

OpenStudy (anonymous):

so it is log b 3=1.099

OpenStudy (anonymous):

wrong answer